A skydriver jump out of a plane traveling 10,000ft above the ground how high is she after 10 seconds? After 15 seconds
Answer by macston(5194)

t=time falling, in seconds
.
Distance fallen=(1/2)(acceleration of gravity)(time in seconds)^2
Distance=1/2(32ft/s^2)(t^2)
For 10 seconds:
Distance fallen=1/2(32ft/s^2)(10s)^2=(16ft/s^2)(100s^2)=1600 feet
Height=10000ft-1600ft=8400 feet
ANSWER 1: She is 8400 feet high after 10 seconds.
For 15 seconds:
Distance fallen=1/2(32ft/s^2)(15s)^2=(16ft/s^2)(225s^2)=3600 feet
Height=10000ft-3600ft=6400 feet
ANSWER 2: She is 6400 feet high after 15 seconds.